# Exact calculation of degrees for lattice equations: a singularity approach

### v1

@@ 1. Introduction
It has been more than thirty years since the study of integrability tests in the discrete systems started. Integrability tests are criteria that predict whether or not a given equation is integrable.
The first integrability test for discrete systems in the history is singularity confinement [10]. It was proposed as a discrete analogue of the so-called Painlevé property. Roughly speaking, a discrete equation enters a singularity if it looses information on initial values, and an equation is said to pass the singularity confinement test if the lost information is recovered after a finite number of iterations. However, it is widely known today that singularity confinement is not perfect as an integrability test. Not only the so-called Hietarinta-Viallet equation [18] but a large number of non-integrable systems that pass the singularity confinement test have been discovered and constructed. It is also known that most linearizable equations possess non-confined singularities [36].
@@
which, exactly speaking, are not singularities since the equation does not loose information on initial values. Note that these patterns extend in both sides. Geometrically speaking, each pattern corresponds to the motion of curves
@@
In addition to these two, there are some minor problems. For example, the number of $z^{*}$ that generates a pattern counted with multiplicity should be defined rigorously in the general cases. Moreover, it is not very easy in general to confirm that a singularity pattern is open, which is not a problem of the method but of singularity analysis.
@@
and $\Phi$ depends on all of $B$, $C$ and $D$. Throughout this paper, we always think of the equation as defining the evolution in the north-east direction. We will introduce some conditions on $\Phi$ (see Definitions 3.6 and 3.7) but do not assume other nonessential conditions. For example, we do not assume that the equation can be solved in another direction.
The paper is organized as follows. In §2, we introduce the lattice version of Halburd’s method without rigorous discussion. We will discuss what problems must be solved to make the method rigorous. The main part of this paper is §3 and §4. We will state our main theorems in §3 and give proofs in §4. Since all we need to rigorously calculate degrees are stated in §3, a reader who is not interested in proofs can skip §4. In §5, we analyze several examples to see how to calculate exact degrees with our theory. For a reader who is not familiar with singularity analysis, we discuss how to detect and calculate singularity patterns, too. In Appendix A, we prove Theorem A.1, which states that when considering the degrees for a quad equations, there are only four essential choices of domain. In Appendix B, we recall some definitions and basic properties of substitution and derivative, which play an important role in §4. Appendix C consists of some algebraic lemmas, which we use in §4. In Appendix D, we recall the definition of divisors on $\mathbb{P}^{1}$ and introduce some unusual notations.
@@
  • ($\mathbb{Z}^{2}_{\leq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\leq t;\ n^{\prime}\leq n\}$, $H_{\leq(t,n)}=H\cap\mathbb{Z}^{2}_{\leq(t,n)}$: the past light-cone emanating from $(t,n)$.
  • ($\mathbb{Z}^{2}_{\geq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\geq t;\ n^{\prime}\geq n\}$, $H_{\geq(t,n)}=H\cap\mathbb{Z}^{2}_{\geq(t,n)}$: the future light-cone emanating from $(t,n)$.
@@ 2. Lattice version of Halburd’s method
In this section, we see on an example how to extend Halburd’s method to lattice equations. It should be noted in advance, however, that while the calculations in the example are correct in the end (Example 5.4), some discussion is not rigorous as Example 1.1. After the example, we will discuss what problems exist and how to fix them.
@@
where “reg” means that the leading order with respect to $\varepsilon$ is 00 and the leading coefficient depends on some other initial variable. Let us fix an initial value problem as
@@
The first problem is the definition of singularities in the case of lattice equations. Although singularity confinement for lattice equations was applied in the original paper [10], the search for its rigorous definition is still going on [44, 45, 35]. Moreover, roughly speaking, all the singularities in our settings should have codimension one since all the initial values other than $z$ are taken generic. Therefore, we must start with rigorously defining what is a singularity we consider in this paper.
@@
$\times$$\times$$\times$$\beta_{1}$
$\times$$\times$$\times$$\beta_{2}$
$z^{*}$
@@
$\times$$\times$$\times$$\beta_{1}$
$\times$$\times$$\times$$\beta_{2}$
$z^{*}$
@@
$\times$$\times$$\times$$\times$$\times$$\times$$\times$$\beta_{1}$
$\times$$\times$$\times$$\beta_{2}$
$z^{*}$
@@
The fourth problem is, as mentioned in the above example, how to confirm that there is no possibility for $x_{t,n}(z)$ to become 00 or $\infty$ except for the patterns $\text{I}_{t,n}$ and II. For example, we must at least check that the spontaneous occurrence of $x_{t,n}(z^{*})=\beta\notin\{0,\infty\}$ does not generate 00 or $\infty$ as a value. This is the most difficult part in our theory and we must develope some techniques to resolve this problem.
The only remedy known today for these problems would be the coprimeness property. The coprimeness property was invented as an algebraic reinterpretation of singularity confinement and can deal with global singularity structure [26]. For example, in the case of Hirota’s discrete KdV equation, one can show that each iterate $x_{t,n}$ decomposes into
@@ 3. How to calculate degrees for lattice equations
  • For any $(t,n)\in H$, the intersection between the corresponding past light-cone and the domain $H$, i.e.,
@@
  • For any $(t,n)\in H$, the corresponding future light-cone
@@
As seen in the rest of the paper, this approach makes all definitions and discussion much clearer. A key idea in this paper is to clarify on which initial variables a value $z^{*}$ we substitute for $z$ depends. In fact, some definitions, lemmas and propositions entirely rely on this approach. The definition of a singularity we use in this paper, which is given below, is such an example.
@@
    If $\{\beta_{1},\ldots,\beta_{I}\}\subset S$ and $I<J$, then the degree relations completely determine all the degrees. That is, solving a system of linear equations, one can compute degrees only from the degree relations.
@@
If the starting value of a movable pattern is $\beta$, then the pattern obviously contains $\beta$ as a value. Therefore, $\alpha\in S\setminus\{\beta_{1},\ldots,\beta_{I}\}$ implies that the equation does not have a movable pattern starting with $\alpha$. In short, to calculate degrees in our method, it is essential for an equation to have a $\mathbb{P}^{1}(\mathbb{C})$-value that does not appear as the starting value of any movable pattern.
@@
$z^{*}$
$w$
Figure 5. Situation in Proposition 3.17. If the points marked with $\boxplus$” are not constant singularities, then those marked with “$+$” depend on $w$ and are not constant singularities, either. This procedure can easily be repeated and thus none of the points east to but not higher than the $\boxplus$-wall is a constant singularity. To prove this proposition, it is essential that the points marked with “$+$” lie outside the past light-cone emanating from $(t,n)$.
@@ 4. Proofs
$x_{t,n}$
$z$
Figure 8. Situation in Definition 4.2. The initial variables in the boxed regions, which lie in the past light-cone emanating from $(t,n)$ but are different from $z$, belong to $\mathbb{K}_{t,n}$.
@@
Here, we simply used $x_{t_{0}-1,n_{0}+1}$ and $x_{t_{0}-1,n_{0}}$ since $x_{t_{0}-1,n_{0}+1}(z)$ and $x_{t_{0}-1,n_{0}}(z)$ do not depend on $z=x_{t_{0},n_{0}}$. By Proposition 3.17, $x_{t_{0}-1,n_{0}+1}$ depends on the initial variable $x_{t_{0}-t^{\prime},n_{0}+1}$ where $t^{\prime}$ is the least positive integer satisfying $(t_{0}-t^{\prime},n_{0}+1)\in H_{0}$. Since $\Phi(B,x^{*},D)\notin\mathbb{P}^{1}(\mathbb{C})$, $x_{t_{0},n_{0}+1}(x^{*})$ must depend on at least one of the initial variables $x_{t_{0}-1,n_{0}}$ and $x_{t_{0}-t^{\prime},n_{0}+1}$. On the other hand, if $(t_{0},n_{0}+1)\in H_{0}$, then $x_{t_{0},n_{0}+1}$ itself is an initial variable. In both cases, $x_{t_{0},n_{0}+1}(x^{*})$ is not a constant singularity. Interchanging the $t$- and $n$- axes, one can show that $x_{t_{0}+1,n_{0}}(x^{*})$ is not a constant singularity, either. Since $x_{t_{0},n_{0}+1}(x^{*})$ and $x_{t_{0}+1,n_{0}}(x^{*})$ depend on different initial variables, they are algebraically independent over $\mathbb{C}$. Therefore, using $\Phi(B,C,x^{*})\notin\mathbb{P}^{1}(\mathbb{C})$, we have
@@
Therefore, $F$ is a non-trivial factor. Next, suppose that the relation $\Phi(B,C,D)=x^{*}$ has a non-trivial factor $F=F(B,C,D)\in\mathbb{C}[B,C,D]$. We show that the equation has a movable pattern starting with $x^{*}$. It follows from the $\partial$-factor condition that $F$ depends on both $B$ and $C$ since $F$ appears in the numerator of $\partial_{B}\Phi$ or $\partial_{B}\frac{1}{\Phi}$ (and $\partial_{C}\Phi$ or $\partial_{C}\frac{1}{\Phi}$) as a factor. Take a lattice point $(t,n)\in H\setminus H_{0}$ that satisfies
@@
Let $n^{\prime}$ be the least nonnegative integer satisfying $(t,n-n^{\prime})\in H_{0}$ and let $w=x_{t,n-n^{\prime}}$. We think of $x_{t,n-1}$ as a rational function in $w$, i.e., $x_{t,n-1}=x_{t,n-1}(w)$. Note that $x_{t-1,n}$ and $x_{t-1,n-1}$ are independent of $w$. Consider the algebraic equation
@@ 5. Examples
where $\Phi(B,C,D)$ is a polynomial in $B$, $C$ and $D$. In this case, $\infty$ cannot appear in any movable pattern. Moreover, only the fixed pattern corresponding to $z=\infty$ contains $\infty$ as a value. Therefore, letting $S=\{\infty\}$ and calculating the multiplicity of $x_{t,n}=\infty$ for this fixed pattern, one can compute the degrees. This result does not rely on the form of the equation except that it is defined by a polynomial. Neither the $\partial$-factor condition nor the basic pattern condition is required. Note that computing exact degrees for polynomial mappings by Halburd’s method is already mentioned in the original paper [15].
@@
The second simplest class of examples would be equations with the Laurent property. An equation (or an initial value problem for an equation) is said to have the Laurent property if every iterate can be expressed as a Laurent polynomial in the initial variables [8]. Since the denominator of each iterate is a monomial, $\infty$ cannot appear in any movable pattern and only the fixed patterns corresponding to $z=0,\infty$ can contain $\infty$ as a value. Therefore, letting $S=\{\infty\}$ and calculating the multiplicity of $x_{t,n}=\infty$ for these two fixed patterns, one can compute the degrees. As in the case of polynomial mappings, this result does not rely on the form of the equation, either.
@@
where $a$ and $b$ are non-zero parameters and $k$ is a positive even integer. This equation was discovered by the present author and reported in detail with collaborators as the first ever example of a non-integrable lattice equation that passes the singularity confinement test [28, 25]. Let $z=x_{t_{0},n_{0}}$ be an arbitrary initial variable and let
@@
Directly calculating $\partial_{B}\Phi$, $\partial_{B}\frac{1}{\Phi}$, $\partial_{C}\Phi$ and $\partial_{C}\frac{1}{\Phi}$, one can easily show that $\Phi$ satisfies the $\partial$-factor condition. Looking at the equation carefully, we notice that it is not easy for $x_{t,n}$ to become $\infty$. In fact, the relation $\Phi(B,C,D)=\infty$ can be written as
@@
A direct calculation shows that $x_{t+1,n}(z^{*})$ depends on $w$, too. Interchanging the $t$- and $n$- axes, one can show that $x_{t,n+1}(z^{*})$ depends on some initial variable $w^{\prime}=x_{t-t^{\prime},n+1}$. Since $x_{t+1,n}(z^{*})$ (resp. $x_{t,n+1}(z^{*})$) is independent of $w^{\prime}$ (resp. $w$), $x_{t+1,n}(z^{*})$ and $x_{t,n+1}(z^{*})$ are algebraically independent over $\mathbb{C}$ and can be thought of as if they were initial variables when calculating $x_{t+1,n+1}(z^{*})$ from $x_{t,n+1}(z^{*})$, $x_{t+1,n}(z^{*})$ and $x_{t,n}(z^{*})$. Therefore, a direct calculation leads to $x_{t+1,n+1}(z^{*})=\infty$. Repeating this procedure, one obtains the whole pattern. On the other hand, the fixed pattern corresponding to $z=0$ varies depending on the shape of the domain, especially on whether or not $(t_{0}+1,n_{0})$ and $(t_{0},n_{0}+1)$ belong to $H\setminus H_{0}$, respectively. If neither $(t_{0}+1,n_{0})$ nor $(t_{0},n_{0}+1)$ belongs to $H_{0}$, then the pattern is
@@
Let us see the calculations in Example 2.1 from a point of view of our theory. One can check that all the calculations are justified by our theory. For example, using our theory, even those without any knowledge of singularity analysis can rigorously calculate the fixed pattern corresponding to $z=\infty$. Since the discussion is the same as in Example 5.3, we omit the details here. It should be noted, however, that there is a small difference between the $Z_{t,n}$ in Example 2.1 and the $Z_{t,n}(0)$ defined in §3 and §4. The difference arises in the fixed pattern starting with $x_{1,1}(z^{*})=0$, i.e., corresponding to $z=0$. While $Z_{1,1}$ contains the information on this pattern, $Z_{1,1}(0)$ does not. This pattern behaves as if it were a movable pattern but is labeled fixed by the strict definition of singularity patterns (Definition 3.4). Therefore, with the notation in Theorem 3.12, the degree relations are now written for $(t,n)\in H$ as
@@
However, $f_{1}$ is not divisible by $B-C$, which is a contradiction. Therefore, $f$ is irreducible and thus the equation satisfies the basic pattern condition for $x^{*}$. Next, we consider the case $x^{*}=\infty$. In this case, the relation $\Phi(B,C,D)=\infty$ is

### v2

@@ 1. Introduction
It has been more than thirty years since the study of integrability tests in discrete systems started. Integrability tests are criteria that predict whether or not a given equation is integrable.
The first integrability test for discrete systems in history is singularity confinement [10]. It was proposed as a discrete analogue of the so-called Painlevé property. Roughly speaking, a discrete equation enters a singularity if it loses information on initial values, and an equation is said to pass the singularity confinement test if the lost information is recovered after a finite number of iterations. However, it is widely known today that singularity confinement is not perfect as an integrability test. Not only the so-called Hietarinta-Viallet equation [18] but a large number of non-integrable systems that pass the singularity confinement test have been discovered and constructed. It is also known that most linearizable equations possess non-confined singularities [36].
@@
which, exactly speaking, are not singularities since the equation does not lose information on initial values. Note that these patterns extend to both sides. Geometrically speaking, each pattern corresponds to the motion of curves
@@
First, let us verify all the possibilities where ±$\pm 3a$ or $\infty$ appears as a value. According to the singularity analysis above, we know the following:
@@
In addition to these two, there are some minor problems. For example, the number of $z^{*}$ that generates a pattern counted with multiplicity should be defined rigorously in general cases. Moreover, it is not very easy in general to confirm that a singularity pattern is open, which is not a problem of the method but of singularity analysis.
@@
and $\Phi$ depends on all of $B$, $C$ and $D$. Throughout this paper, we always think of the equation as defining the evolution in the northeast direction. We will introduce some conditions on $\Phi$ (see Definitions 3.6 and 3.7) but do not assume other nonessential conditions. For example, we do not assume that the equation can be solved in another direction.
The paper is organized as follows. In §2, we introduce the lattice version of Halburd’s method without rigorous discussion. We will discuss what problems must be solved to make the method rigorous. The main part of this paper is §3 and §4. We will state our main theorems in §3 and give proofs in §4. Since all we need to rigorously calculate degrees are stated in §3, a reader who is not interested in proofs can skip §4. In §5, we analyze several examples to see how to calculate exact degrees with our theory. For a reader who is not familiar with singularity analysis, we discuss how to detect and calculate singularity patterns, too. In Appendix A, we prove Theorem A.1, which states that when considering the degrees for a quad equation, there are only four essential choices of domain. In Appendix B, we recall some definitions and basic properties of substitution and derivative, which play an important role in §4. Appendix C consists of some algebraic lemmas, which we use in §4. In Appendix D, we recall the definition of divisors on $\mathbb{P}^{1}$ and introduce some unusual notations.
@@
  • ($\mathbb{Z}^{2}_{\leq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\leq t;\ n^{\prime}\leq n\}$, $H_{\leq(t,n)}=H\cap\mathbb{Z}^{2}_{\leq(t,n)}$: the past light cone emanating from $(t,n)$.
  • ($\mathbb{Z}^{2}_{\geq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\geq t;\ n^{\prime}\geq n\}$, $H_{\geq(t,n)}=H\cap\mathbb{Z}^{2}_{\geq(t,n)}$: the future light cone emanating from $(t,n)$.
@@ 2. Lattice version of Halburd’s method
In this section, we see on an example how to extend Halburd’s method to lattice equations. It should be noted in advance, however, that while the calculations in the example are correct in the end (Example 5.4), some discussion is not rigorous, as Example 1.1. After the example, we will discuss what problems exist and how to fix them.
@@
where “reg” means that the leading order with respect to $\varepsilon$ is 00 and the leading coefficient depends on some other initial variable.
Let us fix an initial value problem as
@@
The first problem is the definition of singularities in the case of lattice equations. Although singularity confinement for lattice equations was applied in the original paper [10], the search for its rigorous definition is still going on [44, 45, 35]. Moreover, roughly speaking, all the singularities in our settings should have codimension one since all the initial values other than $z$ are taken generic. Therefore, we must start by rigorously defining what is a singularity we consider in this paper.
@@
$\times$$\times$$\times$$\beta_{1}$$\times$$\times$$\times$$\beta_{2}$$z^{*}$
@@
$\times$$\times$$\times$$\beta_{1}$$\times$$\times$$\times$$\beta_{2}$$z^{*}$
@@
$\times$$\times$$\times$$\times$$\times$$\times$$\times$$\beta_{1}$$\times$$\times$$\times$$\beta_{2}$$z^{*}$
@@
The fourth problem is, as mentioned in the above example, how to confirm that there is no possibility for $x_{t,n}(z)$ to become 00 or $\infty$ except for the patterns $\text{I}_{t,n}$ and II. For example, we must at least check that the spontaneous occurrence of $x_{t,n}(z^{*})=\beta\notin\{0,\infty\}$ does not generate 00 or $\infty$ as a value. This is the most difficult part of our theory and we must develop some techniques to resolve this problem.
The only remedy known today for these problems would be the coprimeness. This property was invented as an algebraic reinterpretation of singularity confinement and can deal with global singularity structure [26]. For example, in the case of Hirota’s discrete KdV equation, one can show that each iterate $x_{t,n}$ decomposes into
@@ 3. How to calculate degrees for lattice equations
  • For any $(t,n)\in H$, the intersection between the corresponding past light cone and the domain $H$, i.e.,
@@
  • For any $(t,n)\in H$, the corresponding future light cone
@@
As seen in the rest of the paper, this approach makes all definitions and discussion much clearer. A key idea in this paper is to clarify on which initial variables a value $z^{*}$ we substitute for $z$ depends. Some definitions, lemmas and propositions entirely rely on this approach. The definition of a singularity we use in this paper, which is given below, is such an example.
@@
    If $\{\beta_{1},\ldots,\beta_{I}\}\subset S$ and $I<J$, then the degree relations completely determine all the degrees. That is, by solving a system of linear equations, one can compute degrees only from the degree relations.
@@
If the starting value of a movable pattern is $\beta$, then the pattern obviously contains $\beta$ as a value. Therefore, $\alpha\in S\setminus\{\beta_{1},\ldots,\beta_{I}\}$ implies that the equation does not have a movable pattern starting with $\alpha$. In short, to calculate degrees in our method, an equation needs to have a $\mathbb{P}^{1}(\mathbb{C})$-value that does not appear as the starting value of any movable pattern.
@@
Figure 5. Situation in Proposition 3.17. If the points marked with $\boxplus$” are not constant singularities, then those marked with “$+$” depend on $w$ and are not constant singularities, either. This procedure can easily be repeated and thus none of the points east to but not higher than the $\boxplus$-wall is a constant singularity. To prove this proposition, it is essential that the points marked with “$+$” lie outside the past light cone emanating from $(t,n)$.
@@ 4. Proofs
$x_{t,n}$$z$
Figure 8. Situation in Definition 4.2. The initial variables in the boxed regions, which lie in the past light cone emanating from $(t,n)$ but are different from $z$, belong to $\mathbb{K}_{t,n}$.
@@
Here, we simply used $x_{t_{0}-1,n_{0}+1}$ and $x_{t_{0}-1,n_{0}}$ since $x_{t_{0}-1,n_{0}+1}(z)$ and $x_{t_{0}-1,n_{0}}(z)$ do not depend on $z=x_{t_{0},n_{0}}$. By Proposition 3.17, $x_{t_{0}-1,n_{0}+1}$ depends on the initial variable $x_{t_{0}-t^{\prime},n_{0}+1}$ where $t^{\prime}$ is the least positive integer satisfying $(t_{0}-t^{\prime},n_{0}+1)\in H_{0}$. Since $\Phi(B,x^{*},D)\notin\mathbb{P}^{1}(\mathbb{C})$, $x_{t_{0},n_{0}+1}(x^{*})$ must depend on at least one of the initial variables $x_{t_{0}-1,n_{0}}$ and $x_{t_{0}-t^{\prime},n_{0}+1}$. On the other hand, if $(t_{0},n_{0}+1)\in H_{0}$, then $x_{t_{0},n_{0}+1}$ itself is an initial variable. In both cases, $x_{t_{0},n_{0}+1}(x^{*})$ is not a constant singularity. Interchanging the $t$- and $n$- axes, one can show that $x_{t_{0}+1,n_{0}}(x^{*})$ is not a constant singularity, either.
Since $x_{t_{0},n_{0}+1}(x^{*})$ and $x_{t_{0}+1,n_{0}}(x^{*})$ depend on different initial variables, they are algebraically independent over $\mathbb{C}$. Therefore, using $\Phi(B,C,x^{*})\notin\mathbb{P}^{1}(\mathbb{C})$, we have
@@
Therefore, $F$ is a non-trivial factor.
Next, suppose that the relation $\Phi(B,C,D)=x^{*}$ has a non-trivial factor $F=F(B,C,D)\in\mathbb{C}[B,C,D]$. We show that the equation has a movable pattern starting with $x^{*}$. It follows from the $\partial$-factor condition that $F$ depends on both $B$ and $C$ since $F$ appears in the numerator of $\partial_{B}\Phi$ or $\partial_{B}\frac{1}{\Phi}$ (and $\partial_{C}\Phi$ or $\partial_{C}\frac{1}{\Phi}$) as a factor. Take a lattice point $(t,n)\in H\setminus H_{0}$ that satisfies
@@
Let $n^{\prime}$ be the least nonnegative integer satisfying $(t,n-n^{\prime})\in H_{0}$ and let $w=x_{t,n-n^{\prime}}$. We think of $x_{t,n-1}$ as a rational function in $w$, i.e., $x_{t,n-1}=x_{t,n-1}(w)$. Note that $x_{t-1,n}$ and $x_{t-1,n-1}$ are independent of $w$.
Consider the algebraic equation
@@ 5. Examples
where $\Phi(B,C,D)$ is a polynomial in $B$, $C$ and $D$. In this case, $\infty$ cannot appear in any movable pattern. Moreover, only the fixed pattern corresponding to $z=\infty$ contains $\infty$ as a value. Therefore, by letting $S=\{\infty\}$ and calculating the multiplicity of $x_{t,n}=\infty$ for this fixed pattern, one can compute the degrees. This result does not rely on the form of the equation except that it is defined by a polynomial. Neither the $\partial$-factor condition nor the basic pattern condition is required. Note that computing exact degrees for polynomial mappings by Halburd’s method is already mentioned in the original paper [15].
@@
The second simplest class of examples would be equations with the Laurent property. An equation (or an initial value problem for an equation) is said to have the Laurent property if every iterate can be expressed as a Laurent polynomial in the initial variables [8]. Since the denominator of each iterate is a monomial, $\infty$ cannot appear in any movable pattern and only the fixed patterns corresponding to $z=0,\infty$ can contain $\infty$ as a value. Therefore, by letting $S=\{\infty\}$ and calculating the multiplicity of $x_{t,n}=\infty$ for these two fixed patterns, one can compute the degrees. As in the case of polynomial mappings, this result does not rely on the form of the equation, either.
@@
where $a$ and $b$ are non-zero parameters and $k$ is a positive even integer. This equation was discovered by the present author and reported in detail with collaborators as the first-ever example of a non-integrable lattice equation that passes the singularity confinement test [28, 25]. Let $z=x_{t_{0},n_{0}}$ be an arbitrary initial variable and let
@@
Directly calculating $\partial_{B}\Phi$, $\partial_{B}\frac{1}{\Phi}$, $\partial_{C}\Phi$ and $\partial_{C}\frac{1}{\Phi}$, one can easily show that $\Phi$ satisfies the $\partial$-factor condition.
Looking at the equation carefully, we notice that it is not easy for $x_{t,n}$ to become $\infty$. In fact, the relation $\Phi(B,C,D)=\infty$ can be written as
@@
A direct calculation shows that $x_{t+1,n}(z^{*})$ depends on $w$, too. Interchanging the $t$- and $n$- axes, one can show that $x_{t,n+1}(z^{*})$ depends on some initial variable $w^{\prime}=x_{t-t^{\prime},n+1}$. Since $x_{t+1,n}(z^{*})$ (resp. $x_{t,n+1}(z^{*})$) is independent of $w^{\prime}$ (resp. $w$), $x_{t+1,n}(z^{*})$ and $x_{t,n+1}(z^{*})$ are algebraically independent over $\mathbb{C}$ and can be thought of as if they were initial variables when calculating $x_{t+1,n+1}(z^{*})$ from $x_{t,n+1}(z^{*})$, $x_{t+1,n}(z^{*})$ and $x_{t,n}(z^{*})$. Therefore, a direct calculation leads to $x_{t+1,n+1}(z^{*})=\infty$. Repeating this procedure, one obtains the whole pattern.
On the other hand, the fixed pattern corresponding to $z=0$ varies depending on the shape of the domain, especially on whether or not $(t_{0}+1,n_{0})$ and $(t_{0},n_{0}+1)$ belong to $H\setminus H_{0}$, respectively. If neither $(t_{0}+1,n_{0})$ nor $(t_{0},n_{0}+1)$ belongs to $H_{0}$, then the pattern is
@@
Let us see the calculations in Example 2.1 from a point of view of our theory. One can check that all the calculations are justified by our theory. For example, using our theory, even those without any knowledge of singularity analysis can rigorously calculate the fixed pattern corresponding to $z=\infty$. Since the discussion is the same as in Example 5.3, we omit the details here.
It should be noted, however, that there is a small difference between the $Z_{t,n}$ in Example 2.1 and the $Z_{t,n}(0)$ defined in §3 and §4. The difference arises in the fixed pattern starting with $x_{1,1}(z^{*})=0$, i.e., corresponding to $z=0$. While $Z_{1,1}$ contains the information on this pattern, $Z_{1,1}(0)$ does not. This pattern behaves as if it were a movable pattern but is labeled fixed by the strict definition of singularity patterns (Definition 3.4). Therefore, with the notation in Theorem 3.12, the degree relations are now written for $(t,n)\in H$ as
@@
However, $f_{1}$ is not divisible by $B-C$, which is a contradiction. Therefore, $f$ is irreducible and thus the equation satisfies the basic pattern condition for $x^{*}$.
Next, we consider the case $x^{*}=\infty$. In this case, the relation $\Phi(B,C,D)=\infty$ is